OpenStudy (calculusxy):

MEDAL Two-fifths of what number subtracted from 18 is equal to 4 more than two-thirds of 6? F. 4 G. 12 H. 25 J. 32

OpenStudy (calculusxy):

@Mertsj

OpenStudy (calculusxy):

@satellite73

OpenStudy (mertsj):

\[18-\frac{2}{5}x=4+\frac{2}{3}(6)\]

OpenStudy (calculusxy):

\[18-\frac{ 2 }{ 5 }x=4+4\]

OpenStudy (mertsj):

yep

OpenStudy (calculusxy):

\[18-\frac{ 2 }{ 5 }x=8\]

OpenStudy (calculusxy):

Now do I add 18 to 8?

OpenStudy (mertsj):

1. Subtract 18 from both sides 2. Multiply both sides by -5/2

OpenStudy (calculusxy):

I subtract 18 because it's positive 18 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What you do to one side you must do the other, isolate x.